OpenStudy (anonymous):

2(3x – 4) = 3x + 7 can some one solve pleasee

OpenStudy (anonymous):

its just like untying a knot. multiply out the left bracket first 6x - 8 = 3x +7 get all the x terms together 6x -3x = 7 +8 3x = 15 divide by 3 to win x = 5